Suppose that your firm is trying to decide between two machines that will do the same job. Machine A costs $50,000, will last for ten years and will require operating costs of $5,000 per year. At the end of ten years it will be scrapped for $10,000. Machine B costs $60,000, will last for seven years and will require operating costs of $6,000 per year. At the end of seven years it will be scrapped for $5,000. Which of the following statement is correct? (discount rate is 10 percent)

Question 11 options:

B is a better machine because it has a larger NPV.

A is a better machine because it has a larger NPV.

B is a better machine because it has a larger Equivalent Annual Cost.

B is a better machine because it has a smaller Equivalent Annual Cost.

A is a better machine because it has a smaller Equivalent Annual Cost.
Answer #1

A is a better machine because it has a smaller Equivalent Annual Cost.

EAC of A=(50000-10000/1.1^10)*10%/(1-1/1.1^10)+5000=12509.8158

EAC of B=(60000-5000/1.1^7)*10%/(1-1/1.1^7)+6000=17797.3025